OpenStudy (anonymous):

how to solve for p: 3/5 (p-3) = -4

OpenStudy (anonymous):

multiply p-3 on both sides then try to solve

OpenStudy (anonymous):

times both sides by 5 ... -20=3(p-3) . open up the brackets .. -20=3p+9 u get -29/3=p

OpenStudy (anonymous):

no no look at the question this 3 is not multiplied with p-3, it is divided by 3 @wannawarner

OpenStudy (anonymous):

I don't know if either of you guys are right haha. I think you're supposed to subtract 3/4 from both sides.

OpenStudy (anonymous):

@wannawarner @03453660

OpenStudy (anonymous):

**3/5

OpenStudy (anonymous):

so its 3/5 times (p-3) yes ??

OpenStudy (anonymous):

what do you mean @wannawarner

OpenStudy (anonymous):

like on the LHS of the equation

OpenStudy (anonymous):

I got -23/5 for the answer, I'm going to check it real quick @wannawarner @03453660

OpenStudy (anonymous):

ok i do this question 3/5(p-3)=-4 multiply both sides by p-3 3/5=-4(p-3) 3/5=-4p+12 (3/5)-12=4p p= ((3/5)-12)/4

OpenStudy (anonymous):

actually i got -69/5 .....

OpenStudy (anonymous):

One second

OpenStudy (anonymous):

isn,t the ans =-57/20

OpenStudy (anonymous):

yes it is if its the way he says the equation is formated

OpenStudy (anonymous):

ok let him check the ans @wannawarner

OpenStudy (anonymous):

I got p= -8/5.

OpenStudy (anonymous):

can you do the steps

OpenStudy (anonymous):

I mean -11/3

OpenStudy (anonymous):

@03453660 @wannawarner Thanks for your time and help!

OpenStudy (anonymous):

@03453660 Do you still want me to do the steps?

OpenStudy (anonymous):

no no its ok you got it correct

OpenStudy (anonymous):

alright. thanks again.